Using substitution.
X=3y
3x-5y=12

Answer by Simnepi(216)

x = 3y this is equation 1
3x - 5y = 12 this is equation 2
so putting equation 1 into equation 2 gives
3 X 3y - 5y = 12
simplifying gives
9y - 5y = 12
so
4y = 12
therefore
y = 3
substitute this into equation 1
x = 3 X 3
so
x = 9